Giga Manufacturing Co. manufactures 1 GB flash drives (jump drives). Price and cost data for a relevant range extending to 500,000 units per month are as follows:

Sales price per unit:

(Current monthly sales volume is 400,000 units)                       $20.00

Variable costs per unit:

Direct materials                                                                               4.00

Direct labor                                                                                     6.00

Variable manufacturing overhead                                                  2.00

Variable selling and administrative expenses                                 2.00

Monthly fixed expenses:

Fixed manufacturing overhead                                               $1,600,000

Fixed selling and administrative expenses                             $1,200,000

Required:

  1. What would the company’s monthly operating income be if it sold 400,000 units?
  2. Management is currently in contract negotiations with the labor union. If the negotiations fail, direct labor costs will increase by 15% and fixed costs will increase by $400,000 per month. If these costs increase, how many units will the company have to sell each month to break even?
  3. Suppose Giga adds a second line of flash drives (2 GB rather than 1 GB). A package of the 2 GB flash drives will sell for $30 and have variable cost per unit of $20 per unit. The expected sales mix is five of the smaller flash drives (1 GB) for every two larger flash drive (2 GB). Given this sales mix, how many of each type of flash drive will Bytes need to sell to reach its target pre-tax monthly profit of $400,000? Fixed costs will remain the same. Giga’s tax rate is 25%.
Answer #1

a)operating income=sales-variable costs-fixed costs
=(400000*20)-(400000*(4+6+2+2))-(1600000+1200000)=-400,000
so it is net loss
b)breakeven units= fixed costs/(sales-var expenses)
=(1600000+1200000+400000)/(20-4-(6*1.15)-2-2)
=627451 units
c)let units sold for 1gb X and for 2 GB is y
5x=2y and means y=2.5x
contrinbution margin of X=20-(4+6+2+2)=6
contribution margin of Y=20-20=10
profit=sales-variabel cost-fixed expenses
400000=6x+(10*2.5x)-(1600000+1200000)
x=3200000/31
=103226
Y=258065 units